# On the Hardness of Approximation of the Fair $k$-Center Problem

### v1

@@ 3.2 One-per-group regime
Any fair $k$-center with two groups instance can be reduced, in polynomial time, into an equivalent instance of the one-per-group fair $k$-center problem.
@@
    $k$ disjoint groups. Create $r_{1}$ disjoint copies of $G_{1}$ denoted $H^{1}_{1}$, $H^{1}_{2}$, …, $H^{1}_{r_{1}}$, and $r_{2}$ disjoint copies of $G_{2}$ denoted $H^{2}_{1},H^{2}_{2},\ldots,H^{2}_{r_{2}}$. Let $\mathbb{H}=\{H^{1}_{1},\dots,H^{1}_{r_{1}},H^{2}_{1},\dots,H^{2}_{r_{2}}\}$ be the family consisting of these $k=r_{1}+r_{2}$ groups. The requirement vector is $\vec{1}_{k}$, i.e., exactly one center must be selected from each group in $\mathbb{H}$.
@@
$\texttt{cost}_{I}({\widecheck{S}})$ $\leq\alpha\cdot\textsf{\small OPT}(I).$
@@
To conclude the theorem, for the sake of contradiction assume that there exists a polynomial-time $(3-\epsilon)$-approximation algorithm for the one-per-group fair $k$-center problem. By 8, this would imply a polynomial-time $(3-\epsilon)$-approximation algorithm for the fair $k$-center problem with two groups, contradicting Theorem 1. Therefore, unless $\mathsf{P}=\mathsf{NP}$, no polynomial-time $(3-\epsilon)$-approximation algorithm exists for the one-per-group fair $k$-center problem. ∎
@@ 4 Further related work
Our work builds upon broad literature on (fair) clustering and the computational complexity of clustering problems. Since a comprehensive review is beyond the scope of our work, we focus on the results most closely related to our contributions and refer an interested reader to the surveys on (fair) clustering [20, 5].
@@
Clustering is a fundamental problem in computer science, which has been extensively studied in both from theoretical and applied viewpoint [19, 31]. Among many clustering objectives, two well known center-based clustering formulations are $k$-center and $k$-supplier, both of which aim to minimize the maximum distance from each client to its nearest open center. The key distinction lies in the set of admissible centers: in $k$-center, any data point can be chosen as a center, whereas, in $k$-supplier centers must be chosen from a designated subset of data points, referred as facilities [13, 16, 17, 27, 14]. Both problems are $\mathsf{NP}$-hard [31], yet they admit polynomial-time constant-factor approximation algorithms: $k$-center has a $2$-approximation [13, Theorem 2.2], and $k$-supplier has a $3$-approximation [17, Theorem 5]. Moreover, these guarantees are essentially best possible under standard complexity theory assumptions: unless $\mathsf{P}=\mathsf{NP}$, for any $\epsilon>0$, there exists no polynomial-time $(2-\epsilon)$-approximation algorithm for $k$-center [13, Theorem 4.3] and no polynomial-time $(3-\epsilon)$-approximation algorithm for $k$-supplier [17, Theorem 6]. Furthermore, many variants of $k$-center (and $k$-supplier) remain $\mathsf{NP}$-hard, and their inapproximability barriers often persist even under additional constraints. A variant most closely related to our work is $k$-center with costs, for which Chuzhoy et al. [7][Theorem 4.1] prove $(3-\epsilon)$ inapproximability for every $\epsilon>0$, assuming $\mathsf{P}\neq\mathsf{NP}$. This hardness result further extends to the closely related $k$-center problem with forbidden centers, where a subset of points are forbidden to be chosen as centers [1]. Our hardness of approximation results are established via a reduction from $k$-center with forbidden centers.
@@
Algorithmic fairness has gained considerable attention in light of mounting evidence that optimization and learning systems, when designed without accounting for sensitive attributes and structural disparities in the data, can amplify or even create discriminatory outcomes [22, 2]. This has motivated a broader effort to revisit classical combinatorial optimization problems through the lens of fairness, by incorporating additional constraints that in turn require new algorithmic ideas across a range of settings [6, 26, 25, 23, 11, 4, 9, 32]. Within this broader agenda, fairness in clustering has emerged as a particularly active area, driven by the role of clustering in unsupervised decision-making and data summarization. Many notions of fairness have been proposed—capturing proportional representation within clusters or explicit constraints on which representatives are selected—leading to a rich algorithmic literature [12, 15, 24, 11]. In this work, we focus on cluster-center fairness: points are associated with demographic attributes forming groups and fairness constraints are imposed by restricting the number of centers chosen from each group, while optimizing standard objectives such as $k$-median, $k$-means, $k$-center, and $k$-supplier. Existing formulations capture several types of center-selection constraints, including exact requirements [21, 23], lower-bound constraints [30, 29], upper-bound constraints [15, 24, 3], and combined upper-and-lower bound constraints [18, 28, 11]. Our focus in this paper is the fair $k$-center problem with exact constraints.
@@
Kleindessner et al. [23] introduced the fair $k$-center problem in the context of data summarization, and gave a $3\cdot 2^{t-1}$-approximation running in $\mathsf{O}(nkt^{2}+kt^{4})$ time. This was was later improved to a factor-$3$ approximation with running time $\mathsf{O}(nk+n\sqrt{k}\log k)$ by Jones et al. [21]. More recently, Gadekar et al. [10] gave a tight $3$-approximation for fair $k$-supplier in $\mathsf{O}((kn+k^{2}\sqrt{k})\log n\log k)$ time, matching the approximation factor of unconstrained $k$-supplier. All these results assume the standard setting in which the groups form a disjoint partition on the data points. When groups are allowed to overlap, the problem becomes substantially more difficult: in general, even deciding feasibility—i.e., whether there exists any size-$k$ subset of data points that satisfies the fairness requirements for all groups—is $\mathsf{NP}$-hard, which rules out any meaningful approximation guarantees for the clustering objective. Consequently, algorithmic work on intersecting groups typically focuses on parameterized regimes, for example, assuming that the number of centers and/or the number of groups is small or otherwise bounded, under which feasibility and optimization becomes tractable [10]. In a recent work, Gadekar and Thejaswi [11] showed that for certain fair clustering formulations, inapproximability can persist even when feasibility is trivial (i.e., when the fairness requirements satisfiability admits a polynomial-time algorithm). While their result is stated for $k$-median and $k$-means objectives, it suggests a broader phenomenon that may also arise for $k$-center and $k$-supplier. To the best of our knowledge, prior to our work there were no hardness-of-approximation results known for fair $k$-center problem when the groups are disjoint—particularly in the non-degenerate regime where at least one center must be chosen from each group.
@@ 5 Conclusions
Acknowledgments. We thank Aristides Gionis, Bruno Ordozgoiti, and Ameet Gadekar for insightful discussions on the computational complexity of fair clustering over the past several years, which have been an important inspiration for this work and helped to bring closure to the (in)approximability of this problem.

### v2

@@ 3.2 One-per-group regime
Any fair $k$-center with two groups instance can be reduced, in polynomial time, into an equivalent instance of one-per-group fair $k$-center.
@@
    $k$ disjoint groups. Create $r_{1}$ disjoint copies of $G_{1}$ denoted $\{H^{1}_{1}$, $H^{1}_{2}$, …, $H^{1}_{r_{1}}\}$, and $r_{2}$ disjoint copies of $G_{2}$ denoted $\{H^{2}_{1},H^{2}_{2},\ldots,H^{2}_{r_{2}}\}$. Let $\mathbb{H}=\{H^{1}_{1},\dots,H^{1}_{r_{1}},H^{2}_{1},\dots,H^{2}_{r_{2}}\}$ be the family consisting of these $k=r_{1}+r_{2}$ groups. The requirement vector is $\vec{r}=\vec{1}_{k}$, i.e., exactly one center must be selected from each group in $\mathbb{H}$.
@@
$\texttt{cost}_{I}({\widecheck{S}})$ $\leq\alpha\cdot\textsf{\small OPT}(I),$
@@
To conclude, for the sake of contradiction let us assume that there exists a polynomial-time $(3-\epsilon)$-approximation algorithm for the one-per-group fair $k$-center problem. By 8, this would imply a polynomial-time $(3-\epsilon)$-approximation algorithm for the fair $k$-center problem with two groups, contradicting Theorem 1. Therefore, unless $\mathsf{P}=\mathsf{NP}$, no polynomial-time $(3-\epsilon)$-approximation algorithm exists for the one-per-group fair $k$-center problem. ∎
@@ 4 Further Related Work
Our work builds upon broad literature on (fair) clustering and the computational complexity of clustering problems. Since a comprehensive review is beyond the scope of our work, we focus on the results most closely related to our contributions and refer an interested reader to surveys on (fair) clustering [21, 6].
@@
Clustering is a fundamental problem in computer science, which has been extensively studied in both from theoretical and applied viewpoint [20, 32]. Among many clustering objectives, two well known center-based clustering formulations are $k$-center and $k$-supplier, both of which aim to minimize the maximum distance from each client to its nearest center. The key distinction lies in the set of admissible centers: in $k$-center, any data point can be chosen as a center, whereas, in $k$-supplier centers must be chosen from a designated subset of data points, referred as suppliers [14, 17, 18, 28, 15]. Both problems are $\mathsf{NP}$-hard [32], yet they admit polynomial-time constant-factor approximation algorithms: $k$-center has a $2$-approximation [14, Theorem 2.2], and $k$-supplier has a $3$-approximation [18, Theorem 5]. Moreover, these guarantees are essentially best possible under standard complexity theory assumptions: unless $\mathsf{P}=\mathsf{NP}$, for any $\epsilon>0$, there exists no polynomial-time $(2-\epsilon)$-approximation algorithm for $k$-center [14, Theorem 4.3] and no polynomial-time $(3-\epsilon)$-approximation algorithm for $k$-supplier [18, Theorem 6]. Furthermore, many variants of $k$-center (and $k$-supplier) remain $\mathsf{NP}$-hard, and their inapproximability barriers often persist even under additional constraints. A variant most closely related to our work is $k$-center with costs, for which Chuzhoy et al. [8][Theorem 4.1] prove $(3-\epsilon)$ inapproximability for every $\epsilon>0$, assuming $\mathsf{P}\neq\mathsf{NP}$. This hardness result further extends to the closely related $k$-center problem with forbidden centers, where a subset of points are forbidden to be chosen as centers [1]. Our hardness of approximation results are established via a reduction from $k$-center with forbidden centers.
@@
Algorithmic fairness has gained considerable attention in light of mounting evidence that optimization and learning systems, when designed without accounting for sensitive attributes and structural disparities in the data, can amplify or even create discriminatory outcomes [23, 2]. This has motivated a broader effort to revisit classical combinatorial optimization problems through the lens of fairness, by incorporating additional constraints that in turn require new algorithmic ideas across a range of settings [7, 27, 26, 24, 12, 5, 10, 33]. Within this broader agenda, fairness in clustering has emerged as a particularly active area, driven by the role of clustering in unsupervised machine learning and data summarization. Many notions of fairness have been proposed—capturing proportional representation within clusters or explicit constraints on which representatives are selected—leading to a rich algorithmic literature [13, 16, 25, 12]. In this work, we focus on cluster-center fairness: points are associated with demographic attributes forming groups and fairness constraints are imposed by restricting the number of centers chosen from each group, while optimizing standard objectives such as $k$-median, $k$-means, $k$-center, and $k$-supplier. Existing formulations capture several types of center-selection constraints, including exact requirements [22, 24], lower-bound constraints [31, 30], upper-bound constraints [16, 25, 4], and combined upper-and-lower bound constraints [19, 29, 12]. Our focus in this paper is the fair $k$-center problem with exact constraints.
@@
Kleindessner et al. [24] introduced the fair $k$-center problem in the context of data summarization, and gave a $(3\cdot 2^{t-1})$-approximation in $\mathsf{O}(nkt^{2}+kt^{4})$ time. This was was later improved to a factor-$3$ approximation with running time $\mathsf{O}(nk+n\sqrt{k}\log k)$ by Jones et al. [22]. More recently, Gadekar et al. [11] gave a tight $3$-approximation for fair $k$-supplier in $\mathsf{O}((kn+k^{2}\sqrt{k})\log n\log k)$ time, matching the approximation factor of unconstrained $k$-supplier. All these results assume a setting in which the groups form a disjoint partition on the data points. When groups are allowed to overlap, the problem becomes substantially more difficult: in general, even deciding feasibility (i.e., whether there exists any size-$k$ subset of data points that satisfies the fairness requirements for all groups) is $\mathsf{NP}$-hard, which rules out any meaningful approximation guarantees for the clustering objective. Consequently, algorithmic work on intersecting groups typically focuses on parameterized regimes, for example, assuming that the number of centers and/or the number of groups is small or otherwise bounded, under which feasibility and optimization becomes tractable [11, 30, 29, 34]. In a recent work, Gadekar and Thejaswi [12] showed that for certain fair clustering formulations with intersecting groups, inapproximability can persist even when feasibility is trivial (i.e., when satisfying the fairness requirements is decidable in polynomial time). While their results are stated for fair range $k$-median (and $k$-means), it suggests a broader phenomenon that may also arise for fair $k$-center and fair $k$-supplier.
To the best of our knowledge, prior to our work there were no hardness of approximation results known for the fair $k$-center problem with disjoint groups, particularly in the non-degenerate regime where at least one center must be chosen from each group.
@@ 5 Conclusions
Acknowledgments. We thank Aristides Gionis, Bruno Ordozgoiti, and Ameet Gadekar for insightful discussions on the computational complexity of fair clustering over the past several years, which have been an important inspiration for this work.